OpenStudy (anonymous):

a triangle has side lengths 10,15 and 7 is the triangle acute,obtuse,or right show your work and explain please tell me how to get the answer im terrible at geometry

OpenStudy (blurbendy):

Use these rules: 1) a^2 + b^2 = c^2 => right triangle (2) a^2 + b^2 > c^2 => acute-angled (3) a^2 + b^2 < c^2 => obtuse-angled im guessing a = 10 b = 15 and c = 7 but it's hard to say without a picture...

OpenStudy (anonymous):

thye didnt give me a pic

OpenStudy (anonymous):

ok thanks what would you say it is

OpenStudy (anonymous):

well i keep getting different answers i am so confused

OpenStudy (anonymous):

325

OpenStudy (blurbendy):

Actually, hold on, I explained this wrong The two shortest sides are going to be a and b

OpenStudy (blurbendy):

so, what is 10^2 + 7^2 ?

OpenStudy (anonymous):

149

OpenStudy (blurbendy):

Good, and 15^2 = 225 So, 149 < 225 According to our rules, what's the answer?

OpenStudy (anonymous):

i have no idea

OpenStudy (blurbendy):

1) a^2 + b^2 = c^2 => right triangle (2) a^2 + b^2 > c^2 => acute-angled (3) a^2 + b^2 < c^2 => obtuse-angled

OpenStudy (blurbendy):

is 149 = 225?

OpenStudy (anonymous):

no its less would it be acute

OpenStudy (blurbendy):

It is less than, but the less than symbol is "<" not ">", so it's obtuse.

OpenStudy (anonymous):

oops yeah i knew that hehe what would put as th explanation

OpenStudy (blurbendy):

I'm not sure what you mean.

OpenStudy (anonymous):

well i have to explain why thats the answer

OpenStudy (blurbendy):

because the pythagorean theorem says so. If the sum of the two smaller sides squared are less than the hypotenuse squared, then the triangle must be obtuse.

OpenStudy (anonymous):

oh ok thank you i gave you a medal

OpenStudy (blurbendy):

no problem, any time. and thanks!

OpenStudy (anonymous):

your welcome thank you